OpenStudy (anonymous):

tickets are for admission to a museum are $8 for adults, $5 for children and $6 for seniors. what algebraic expression models the total # of dollars cllected?

Parth (parthkohli):

Let \(d = \text{number of dollars collected}\) there it is!

OpenStudy (anonymous):

if 20 adulf tickets, 16 children tickets and 10 senior tickets are sold one morning, how much money is collected in all??

OpenStudy (zzr0ck3r):

if it where just adults I would say d = 8*A where d = dollars made

OpenStudy (anonymous):

i think this one is just looking for answers zz

Parth (parthkohli):

You can collect \(8 \times \text{number of adults}\) from adults. \(5\times {\text{number of children}}\) from children. \(6 \times \text{number of seniors}\) from seniors. The total number of dollars collected will be the sum of all the money collected from all the three categories given.

OpenStudy (anonymous):

i dont know what do for the first one:c

OpenStudy (anonymous):

help????:/